Status Report

ISS On-Orbit Status 14 Aug 2002

By SpaceRef Editor
August 14, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Crew day again began last night at 8:00 pm EDT, ending this morning at 11:30 am.

The ISS residents continued preparing for the spacewalk on Friday morning.  Today’s focus was on Orlan suit fit checks, suited exercise, and communications checks.  [After CDR Korzun set up the CT (communications and tracking) configuration for EVA, the crew first checked out the Orlan systems for proper functioning.  Wearing their biomedical instrumentation belts, they then tested biomedical comm for vital signs monitoring.  After donning equipment, Korzun and Whitson ingressed the suits through their open backs and then closed the backpacks.  During suited exercise, the Orlans and their BSS controls were checked out again, tested for leaks and subjected to fit checks at reduced suit pressure.  Assisted by Sergey Treschev, Valery and Peggy then tried out their ability to move about in the Orlans and evaluated their spatial orientation and positioning in the DC-1 while testing BITS telemetry connections for the EVA.]

During yesterday’s comm checks with the ORLAN suits, the CDR’s primary suit radio produced unexpected noisy tones when Korzun removed the hardline and went to RF (radio frequency) “Korona” mode.  He was able to continue transmitting and receiving communications, but the tones are extremely annoying.  Since the problem did not show up with the backup radio, the Russian specialists have proposed that Korzun perform the EVA using the backup as the operational radio if the tones persist.

After the Robotics checkout ops yesterday, MCC-H powered down the MSS (mobile service system) during crew sleep.  Early in the day, FE-1 Whitson then disconnected the UOP (utility operations panel) bypass cable to the Lab RWS DCP (robotic workstation/display & control panel), which she had hooked up yesterday.

Yesterday’s ground-controlled checkouts of the robotics MBS (mobile base system) mast camera and MCAS (MBS common attachment system) capture latch assembly (CLA) were completed as planned.  All checkouts were nominal, with two minor exception.  [(1) Right before power-down of the MCAS, a commanded BIT (built-in test) caused the MSS to unexpectedly safe itself (most likely because a position indication set internally to zero when the BIT tests are run is interpreted by the software as inadvertent motion), and (2) slight orange/brown contamination appears to be on the MBS mast camera lens, similar to what has been seen on the robotarm’s cameras.]  

FE-1 Whitson performed a controlled power-down of the VOA (volatile organics analyzer) instrument, prior to MCC-H reactivating it remotely.  This sometimes repeated power-cycling is required to ensure adequate nitrogen flow to VOA.  Later in the day, after warming up, the VOA was ground-commanded to  perform a calibration run of approximately 3.5 hours, followed by a sampling run for another 3.5 hours.

Peggy Whitson was also requested to power cycle the IV-CPDS (intravehicular charged particle directional spectrometer), which was apparently hung up.

On the ADVASC (advanced astroculture), Peggy exchanged/renewed the fluid in the nutrient reservoir, a periodic service task of the plant growth experiment.

She also completed her weekly task of filling out the FFQ (food frequency questionnaire), which keeps a log of her nutritional intake over time.

CDR Korzun and FE-2 Treschev both completed their weekly data collection for the NTXN “Interactions” experiment, entering their perceptions in the mood, group and journal questionnaire on the HRF (human  research facility) PC.  Data are stored in encrypted form on the hard disk, which will later be returned to Earth.

Daily routine tasks were completed by Sergey Treschev (SOSh life support systems maintenance, inventory management system “delta” file preparation) and Peggy Whitson (Lab payloads status checkup).

After several attempts, the crew succeeded yesterday in their troubleshooting to get U.S. EMU (extravehicular mobility unit) batteries to discharge in the Joint Airlock BSA/BCs (battery stowage assembly/battery chargers).  The successful accomplishment of the procedure will make it unnecessary to discharge the batteries in the suit by running the backpack’s fan.  Some more discharge/recharge work needs to be done, but it can wait until after the Orlan EVA-7.

Yesterday’s planned transfer of 277 kg oxidizer (nitrogen tetroxide, N2O4) from the Progress 9P to the ISS using SM compressor #2 was cancelled when the compressor did not start.   [There have been problems with FGB compressor #2 in the past, but for the SM’s compressor #2 this is the first occurrence of an issue.  On the day before (8/12), the transfer of UDMH fuel (unsymmetrical dimethyl hydrazine) using SM compressor #1 was terminated two hours earlier than expected.  Russian specialists think that both problems may be related and that there may be a problem with the compressor-controlling software.  Telemetry data are now being reviewed.]

After MSG/SUBSA (Microgravity Science Glovebox/solidification using a baffle in sealed ampoules) sample operations were halted due to a shattered ampoule in the thermal chamber, POC (Payload Operations Center, Huntsville) has begun developing a plan to get MSG and SUBSA up and running again.  There are no operations planned for either payload until next week, 8/20   For now, MSG will need to remain sealed, and any access to the MSG work volume will have to be via the MSG’s  airlock.

The Elektron oxygen generator system continues to run nominally in 32 Amp mode (three persons), but on its backup pump.  The problems with the low/high sensors on the buffer tank were resolved by end of last week, but then the primary pump shut down.  It is believed that the shutdown was caused by cavitation effects of nitrogen bubbles in the primary pump.  This has been seen before on Mir, where it took about two weeks for the N2 bubbles to settle out of the system.

Today’s targets for CEO (crew earth observations) were Seattle, Wash. (Puget Sound and the urban region were just right of track), Sumatra smoke (crew was to look half right for possible views of Sumatra smoke streaming into the Indian Ocean towards the spacecraft.  Dense smoke from hundreds of fires in Sumatra and Borneo is making news with many SE Asian cities under health alerts), Taiwan Smog (ISS pass was along the axis of the island. Taiwan is one of the smoggiest places in East Asia, with the populated west side of the island often invisible from space), Beijing, China (nadir pass.  Crew to try a mapping series over this large city of 7.1 million [China’s second largest], and to look right for the other city target Tianjin [4.3 million], Beijing’s megaport [and China’s third largest city] on the coast.  Oblique views should have revealed the outlines and direction of movement of the regional smog plume), Cairo, Egypt (nadir pass over this major city of 7.7 million.  One of the fastest growing, the city has gained one million people in seven years.  City margins are changing most rapidly as a result), and Barcelona, Spain (coastal city of 1.5 million under track.  Urban development [mainly for vacationers from northern Europe] has spread along the coasts north and south of Barcelona for tens of km).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:07 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is on Override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 757, temperature (deg C) — 26.7, ppO2 (mmHg) — 149.6, ppCO2 (mmHg) — 3.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 762, temperature (deg C) — 20.4.
  • FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 749.61, temperature (deg C) — 23.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 751.58, temperature (deg C) — 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 751.58, temperature (deg C) — 24.7; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.2
  • PMA-2:  Shell heater temp (deg C) — 18.1

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in directed position (Blind mode, non solar-tracking).
  • SM batteries: Battery #1 is off-line, battery #8 is in Cycle mode; all other batteries (6) are in “Partial charge” mode.
  • FGB batteries:  Battery #5 is off-line; all other batteries (5) are in “Partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM is Off.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line.
  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS:  Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:29 am EDT [= epoch]):

  • Mean altitude — 394.5 km
  • Apogee — 406.5 km
  • Perigee — 382.5 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0017732
  • Orbits per 24-hr. day — 15.57
  • Altitude decrease — 150 m (mean) in last 24 hours
  • Solar Beta Angle — -9.0 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 21311
  • Current Flight Attitude — LVLH (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7.8 deg, roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.